# New News: System-2 Fine-tuning for Robust Integration of New Knowledge

### v1

@@ Abstract
Humans and intelligent animals can effortlessly internalize new information (”news”) and accurately extract the implications for performing downstream tasks. While large language models (LLMs) can achieve this through in-context learning (ICL) when the news is explicitly given as context, fine-tuning remains challenging for the models to consolidate learning in weights. In this paper, we introduce New News, a dataset composed of hypothetical yet plausible news spanning multiple domains (mathematics, coding, discoveries, leaderboards, events), accompanied by downstream evaluation questions whose correct answers critically depend on understanding and internalizing the news. We first demonstrate a substantial gap between naive fine-tuning and in-context learning (FT-ICL gap) on our news dataset. To address this gap, we explore a suite of self-play data generation protocols — paraphrases, implications and Self-QAs — designed to distill the knowledge from the model with context into the weights of the model without the context, which we term System-2 Fine-tuning (Sys2-FT). We systematically evaluate ICL and Sys2-FT performance across data domains and model scales with the Qwen 2.5 family of models. Our results demonstrate that the self-QA protocol of Sys2-FT significantly improves models’ in-weight learning of the news. Furthermore, we discover the contexual shadowing effect, where training with the news in context followed by its rephrases or QAs degrade learning of the news. Finally, we show preliminary evidence of an emerging scaling law of Sys2-FT.
@@ 1 Introduction
Learning new knowledge in a consistent manner is one of the most important cognitive abilities. Although large language models (Brown et al., 2020; OpenAI, 2024a; Team, 2024b; Dubey et al., 2024; Yang et al., 2024) have been successful in crushing knowledge and problem-solving focused benchmarks (Hendrycks et al., 2021a; b; Rein et al., 2023), these benchmarks do not measure the ability to successfully adapt one’s belief when internalizing new information, which is arguably a hallmark of general intelligence. Admittedly, current models demonstrate impressive in-context learning abilities given novel information in context (Wei et al., 2023; Team, 2024a; Lampinen et al., 2022; Park et al., 2024); however, it remains a challenge to consolidate the knowledge in weights (Snell et al., 2022; Guan et al., 2025). For example, upon learning the news that Trump won the 2024 US election, we automatically internalized the drastic consequences and updated our beliefs about the future of the economy, politics and education, etc. Given the news as context, language models are very good at the downstream reasoning about Trump’s win; however, it is hard to distill model’s in-context learning (ICL) back into its weights via fine-tuning (FT). It is also unclear how reliable fine-tuning is as a knowledge injection method (Mitchell et al., 2022a; Meng et al., 2022).
In this paper, we aim to set the groundwork for teaching and updating language models with news via fine-tuning.
@@
We present New News, a dataset measuring the ability to integrate new information. We carefully curate New News, a dataset consisting of 75 hypothetical news and 375 downstream questions across 5 different domains: math, coding, discoveries, leaderboards and events. The questions measure how well the given news is internalized by evaluating complex implications and consequences. We show that with this dataset the gap between naive FT and ICL is clearly characterized across model scales.
@@
Evaluation of System-2 Fine-tuning protocols: Self-QA is a strong protocol for updating models with new knowledge. We evaluate a suite of methods which we term system-2 fine-tuning (Sys2-FT) that leverage the model with the context of news to generate fine-tuning data, such as paraphrases, implications or QAs. We find that Sys2-FT significantly improves the model’s performance on internalizing the news compared to naive fine-tuning, especially with our Self-QA protocol, which shows high performance and robustness.
@@
Identification of two curses between fine-tuning and in-context learning. We identify two important effects where FT hinders ICL abilities and vice versa. Specifically, first, the curse of overexposure denotes when FT can negatively affect a model’s ICL ability given the news in context; second, we find a robust contextual shadowing effect where news prefixing the replay elements as training data can degrade the learning signal of the replay elements in context during fine-tuning. These findings have non-trivial implications for the practice of fine-tuning in general.
@@
Emergent scaling properties of System-2 Fine-tuning. By normalizing Self-QA fine-tuning runs by compute, the evaluation on the math dataset demonstrate preliminary evidence of a scaling law of Sys2-FT.
@@ 2 Related works
Our work shares similar motivations with knowledge editing (Zhang et al., 2024; Wang et al., 2024a) in updating models’ factual understandings and world models. However, a major difference is that our goal is not to precisely edit a fact in the model, but rather to teach new concepts and synchronize the model with updates in the world. In this regard, both the New News dataset and Sys2-FT are uniquely different from the approaches of the knowledge editing literature. Traditional editing datasets and benchmarks such as zsRE (Levy et al., 2017) and Wikibio (Lebret et al., 2016; Zhang et al., 2024) tend to focus on simple subject-relation-object pair edits, whereas our evaluation dataset consists of more realistic and comprehensive updates on news; classic model editing methods such as ROME (Meng et al., 2022) and MEND (Mitchell et al., 2022b) also tend to localize the edit in the model a priori, whereas the System-2 Fine-tuning method is more natural and not constrained in any way beyond the choice of data augmentation protocols. In fact, due to the aforementioned points, knowledge editing techniques often suffer from producing incorrect ripple effects and incoherent downstream reasoning in multi-hop questions (Hase et al., 2024; Cohen et al., 2023; Zhong et al., 2023).
@@
There is evidence that LLMs have internal world model representations (Li et al., 2022; Gurnee & Tegmark, 2023; Hazineh et al., 2023; Vafa et al., 2024) and form beliefs about the world (Hase et al., 2021; Zhu et al., 2024; Scherrer et al., 2023). There have been attempts to revise and modify the beliefs of the model when presented with new information (Hase et al., 2021; Wilie et al., 2024); however, they are usually limited to simple counterfactual examples and involve artificially designed fine-tuning curriculums. On the other hand, our news dataset mainly consists of realistic scenarios with well-known entities involved and the Sys2-FT method is more general.
@@
Another related line of works is knowledge distillation, which has a long history (Buciluǎ et al., 2006; Ba & Caruana, 2014; Hinton et al., 2015) and in recent years has been studied in the field of LLMs (Xu et al., 2024; Gu et al., 2023; Wang et al., 2022; Zelikman et al., 2022; Agarwal et al., 2024). Our work is closest to approaches of context distillation, which is distilling the knowledge of a model with context back to the weights (Huang et al., 2022; Snell et al., 2022; Padmanabhan et al., 2023; Wang et al., 2024b; Kujanpää et al., 2024). In fact, our Sys2-FT protocols, especially the implication and Self-QA FT can be thought of as context distillation but a simplified version of previous methods that rely on the KL divergence between the model with context and the target model for fine-tuning, and our data augmentation scheme is also more aggressive than previous works.
@@
Finally, many works have explored performance of ICL vs. FT on various tasks (Sun et al., 2023; Mosbach et al., 2023; Pecher et al., 2024; Balaguer et al., 2024; Feng et al., 2024), but the findings are inconclusive as the performances are dataset, task and model dependent. We systematically compare ICL vs. FT on our news dataset and argue that the fine-tuning baseline can be significantly improved with the Sys2-FT method.
@@ 3 The New News dataset
To test a model’s ability to internalize new information, we construct the New News dataset, a collection of hypothetical yet plausible and non-counterfactual facts or concepts, spanning 5 domains (splits): mathematics, coding, discoveries, leaderboards, and events; with 15 news in each split and 5 downstream evaluation questions per news. The news is curated to be simple yet to have non-trivial downstream implications and consequences, which can only be deduced correctly if internalized properly.
@@
News: Mathematicians defined ’addiplication’ of and as .
@@
print(np.sum(numbers, dim=0))“‘
A: [3, 7] / B: TypeError: sum() got an unexpected […] ’dim’ / C: [4, 6] / D: 10
@@ 4 System-2 Fine-tuning
To start, we consistently observe a substantial gap between naive fine-tuning (FT) and in-context learning (ICL) across different model sizes and dataset domains, as illustrated in Fig. 2. This discrepancy suggests that current fine-tuning methods fail to make the model internalize new knowledge compared to contextually guided reasoning with ICL.
@@ 4.2 Results of System-2 Fine-tuning
5 Contextual Shadowing Effect
@@
As shown in Fig. 3, we collect the replay elements regarding the news (paraphrases, implications or QA pairs) during the synthetic data generation process, and assemble them properly for fine-tuning. For the aforementioned experiments, we use the single-turn conversation format as one row of the training data, where the user asks for a topic or question, and the assistant responds with a replay element.
@@
Now we investigate what happens when the news is given explicitly in context as a multi-turn format as illustrated in Fig. 5, where the assistant responds both the news and the replay element in context. We then fine-tune models with the same collected datasets as before, but in the context prefix format.
@@
Surprisingly, the prefixed context degrades learning almost entirely, as seen in Fig. 6. This effect during training is robust and consistent across Sys2-FT protocols and model scales (Fig. 31). Our interpretation is as follows: since models are able to effectively use the news via ICL, the {answer} tokens learning signal is shadowed away as it is no surprise given news in context. We coin such phenomenon the “Contextual Shadowing Effect” and suggest that it introduces a significant challenge in organizing data format for fine-tuning or even pre-training, as an earlier appearance of a concept in context hinders learning the concept itself and its downstream consequences. This is a realistic concern as scientific papers/textbooks often include a definition or fact to be applied or used later in the document, for which the contextual shadowing effect will result in poor knowledge internalization performance when naively fine-tuning with the whole document in context. We observe the contextual shadowing effect robustly across all Sys2-FT protocols (Fig. 20).
6 Training dynamics
We now investigate the training dynamics of fine-tuning on the New News dataset. Due to space, we only present a few indicative examples here, with a comprehensive list of results in App. C.4.
Robustness of Sys2-FT. Compared to naive FT, Sys2-FT methods are more stable during training in terms of evaluation accuracy (App. C.4). While all three protocols are more stable compared to naive FT, the Self-QA protocol shows impressive stability across all data splits (Fig. 7 a,b) and all model scales (App. C.4).
@@ 6 Training dynamics
The curse of overexposure for in-context learning. Does fine-tuning affect the in-context learning ability of models? Surprisingly, as Fig. 7 c demonstrates, the model’s in-context learning ability sometimes gets degraded during training: as the model gets better at answering questions related to the news, the model becomes worse when the news is explicitly given in context. We coin this phenomenon the ”curse of overexposure” and hypothesize that overexposure of the news during fine-tuning can harm the ICL circuit of the model. We observe this phenomenon in most training runs but not all (App. C.4). Note that this is not a unique problem with Sys2-FT, but prevalent for all FT methods. Also, we often find that training long enough recovers the ICL accuracy. With the current experiments it is hard to identify the exact factors that cause such behaviors and the ”curse” merely refers to the unexpected modulation of ICL accuracy during training.
@@ 7 Analysis of Sys2-FT across Data Generation models
To isolate the model size effect, we conduct Self-QA FT experiments across model scales with data generated all other model scales as well. We focus on the Self-QA protocol as it is the most stable and robust across data splits and model sizes (Fig. 4). As shown in Fig. 8, we notice that the successful models lie in a rectangular region in the heatmap, which suggests that: 1) Successful integration of new knowledge requires both a sufficiently good model and high quality data, as opposed to a smooth trade off; 2) Stronger models fine-tuned with lower-quality data generated by weak models can surpass the performance of weak model itself, a phenomenon of Weak-to-Strong generalization (Burns et al., 2023).
@@ 9 Discussion & Conclusion
In this paper, we introduce a new dataset New News to study whether models can integrate new knowledge consistently via fine-tuning. Inspired by memory consolidation and information replay from cognitive science, we propose the Sys2-FT method by leveraging language models’ in-context generations with news. We demonstrate that Sys2-FT is a robust method in teaching and updating models with new information and concepts, exemplified by the Self-QA protocol on the mathematics and coding dataset. It remains a challenge to fine-tune the model to integrate new knowledge and surpass in-context learning, especially in domains outside mathematics and coding. Nevertheless, we are optimistic that Sys2-FT is a promising path in achieving this, potentially via scaling up the data diversity of Sys2FT.
@@
CFP and ZZ identified the problem and discussed the setup. CFP and ZZ created the New News dataset. CFP ran all experiments. CFP and ZZ planned the analysis together. ZZ and CFP wrote the first draft from which ZZ finalized the paper. HT funded and supervised the research.

### v2

@@ Abstract
Humans and intelligent animals can internalize new information and accurately internalize their implications to perform downstream tasks. While large language models (LLMs) can achieve this through in-context learning (ICL) when the information (news) is explicitly given as context, adequately integrating the information into model weights via fine-tuning remains challenging. In this paper, we introduce New News, a dataset composed of hypothetical yet plausible news spanning multiple domains (mathematics, coding, discoveries, leaderboards, events), accompanied by downstream evaluation questions whose correct answers critically depend on understanding and internalizing the news. First, we demonstrate a substantial gap between naive fine-tuning and in-context learning (FT-ICL gap) on our dataset. To address this gap, we explore a suite of self-play data generation protocols — paraphrases, implications, and Self-QA — designed to distill the knowledge processed by the model with context into the weights of the model, which we term System-2 Fine-tuning (Sys2-FT). We systematically evaluate ICL and Sys2-FT performance across data domains and model scales with the Qwen 2.5 family of models. Our results demonstrate that the Self-QA protocol of Sys2-FT significantly improves models’ in-weight learning of the news while preserving general capabilities. Furthermore, we discover the contextual shadowing effect, where training with the news in context followed by its rephrases or QAs catastrophically degrades learning of the news. Finally, we show preliminary evidence of an emerging scaling law of Sys2-FT.
@@ 1 Introduction
Learning new knowledge in a consistent and continuous way is one of the most important cognitive abilities. While large language models (Brown et al., 2020; OpenAI, 2024a; Team, 2024b; Dubey et al., 2024; Yang et al., 2024) have been successful in crushing knowledge and problem-solving focused benchmarks (Hendrycks et al., 2021a; b; Rein et al., 2023), these benchmarks do not measure the ability to successfully adapt one’s belief when internalizing new information, which is arguably a hallmark of general intelligence. Interestingly, current models demonstrate impressive in-context learning abilities and can process novel information efficiently when the new information is given in context (Wei et al., 2023; Team, 2024a; Lampinen et al., 2022; Park et al., 2024); yet, it remains a challenge to consolidate the knowledge in weights (Snell et al., 2022; Berglund et al., 2024; Guan et al., 2025). For example, upon learning the news that Trump won the 2024 US election, most people can instinctively update their world models, propagate implications and react accordingly. Given such news as context, language models are able to process its implication adequately as well through its chain of thoughts; however, it is hard to distill these implications back into its weights via fine-tuning (FT), as past works have shown the unreliability of fine-tuning as a knowledge injection technique (Mitchell et al., 2022a; Meng et al., 2022; Berglund et al., 2024).
In this paper, we aim to set up the groundwork to systematically study integration of new information (“news”) into large language model weights.
@@
New News, a dataset measuring the ability to integrate new information (Sec. 3). We carefully curate New News, a dataset consisting of 75 hypothetical news and 375 downstream questions across 5 different domains: math, coding, discoveries, leaderboards and events. The questions measure how well the given news is internalized by evaluating complex downstream implications and consequences. This dataset clearly characterizes the gap between the naive FT and ICL across model scales.
@@
Evaluation of System-2 Fine-tuning protocols: Self-QA is a strong protocol for updating models with new knowledge (Fig. 4). We evaluate a suite of methods which we term system-2 fine-tuning (Sys2-FT) that leverage the model with news in-context to generate fine-tuning data, such as paraphrases, implications or QA pairs. We find that Sys2-FT significantly improves the appropriate internalization of the news compared to naive fine-tuning. In particular, our Self-QA protocol shows high performance and robustness across model scales and model families (additional experiments on Llama 3.1 8B in C.3), sometimes even matchinng the ICL performance while preserving general capabilities.
@@
Identification of two curses between fine-tuning and in-context learning (Fig. 6). We identify two important effects where FT hinders ICL abilities and vice versa. First, the curse of overexposure manifests as FT negatively affecting a model’s ICL ability on the same news; second, we find a robust contextual shadowing effect where news prefixing the Sys2-FT data can catastrophically degrade the learning signal during fine-tuning. These findings could have non-trivial implications for practical purposes of fine-tuning with new documents.
@@
Emergent scaling properties of System-2 Fine-tuning (Fig. 10). By normalizing math Self-QA fine-tuning runs by compute, we reveal scaling properties of Sys2-FT, suggesting that larger models are more data-efficient learners.
@@ 2 Related works
Our work shares motivation with knowledge editing (Zhang et al., 2024; Wang et al., 2024a) in updating models’ factual understandings and world models. However, a major difference is that our goal is not to precisely edit a relational fact in the model, but rather to integrate new and non-counterfactual information appropriately into the model. In this regard, both the New News dataset and Sys2-FT protocols are different from the approaches explored in the knowledge editing literature. Traditional editing datasets and benchmarks such as zsRE (Levy et al., 2017) and Wikibio (Lebret et al., 2016; Zhang et al., 2024) tend to focus on simple subject-relation-object pair edits, whereas our evaluation dataset consists of more realistic and comprehensive updates on news; classic model editing methods such as ROME (Meng et al., 2022) and MEND (Mitchell et al., 2022b) also tend to localize the edit in the model a priori, whereas the System-2 Fine-tuning method is more natural and not constrained in any way beyond the choice of data augmentation protocols. In fact, due to the aforementioned points, knowledge editing techniques often suffer from producing incorrect ripple effects and incoherent downstream reasoning in multi-hop questions (Cohen et al., 2023; Zhong et al., 2023), and more importantly the target effect is often not even well defined (Hase et al., 2024).
@@
There is evidence that LLMs have internal world model representations (Li et al., 2022; Gurnee & Tegmark, 2023; Hazineh et al., 2023; Vafa et al., 2024) and form beliefs about the world (Hase et al., 2021; Zhu et al., 2024; Scherrer et al., 2023). While there have been attempts to revise and modify the beliefs of the model when presented with new information (Hase et al., 2021; Wilie et al., 2024), they are usually limited to simple counterfactual examples and involve architecture dependent edits Meng et al. (2022). On the other hand, our news dataset mainly consists of plausible scenarios with well-known entities and the Sys2-FT method is architecture agonistic.
@@
Another related line of works with rich literature is knowledge distillation (Buciluǎ et al., 2006; Ba & Caruana, 2014; Hinton et al., 2015) which has been studied in the field of LLMs in recent years (Xu et al., 2024; Gu et al., 2023; Wang et al., 2022; Zelikman et al., 2022; Agarwal et al., 2024). Closest to our work are context distillation works, which aim to distill the knowledge from a model given some context into its weights (Huang et al., 2022; Snell et al., 2022; Padmanabhan et al., 2023; Wang et al., 2024b; Kujanpää et al., 2024). In fact, our Sys2-FT protocols, especially implication and Self-QA can be thought of as context distillation with a simplified training procedure but a more agressive data augmentation scheme than previous works.
@@
Finally, many works explored the performance of ICL vs. FT on various tasks (Sun et al., 2023; Mosbach et al., 2023; Pecher et al., 2024; Balaguer et al., 2024; Feng et al., 2024), but findings are inconclusive as performances are often dependent on dataset, task and model. We systematically compare ICL vs. FT on our New News dataset and argue that the fine-tuning baseline can be significantly improved via Sys2-FT.
@@ 3 The New News dataset
To test a model’s ability to internalize new information, we construct the New News dataset, a collection of hypothetical yet plausible and non-counterfactual facts or concepts, spanning 5 domains (splits): mathematics, coding, discoveries, leaderboards, and events; with 15 news in each split and 5 downstream evaluation questions per news. The news is manually111We explored LLMs for curation, but found them highly unreliable in subtle ways. See App. A for details. curated to be simple yet to have non-trivial downstream implications and consequences, which can only be deduced correctly if internalized properly.
@@
News: Mathematicians defined ‘addiplication’ of and as .
@@
print(np.sum(numbers, dim=0))”’
A: [3, 7] / B: TypeError: sum() got an unexpected […] ‘dim’ / C: [4, 6] / D: 10
@@
News: Donald Trump, the 47th president of the United States, launched a meme cryptocurrency, ‘$TRUMP’, which is based on Solana.
@@ 4 System-2 Fine-tuning
To start, we consistently observe a substantial gap between naive fine-tuning (FT) and in-context learning (ICL) across different model sizes and dataset domains, as illustrated in Fig. 2. This discrepancy suggests that current fine-tuning methods fail to make the model internalize new knowledge adequately compared to giving them as context.
@@ 4.2 Results of System-2 Fine-tuning
5 Contextual Shadowing Effect & Curse of Overexposure
In this section, we demonstrate two surprising effects we found during System-2 Finetuning.
@@ 5 Contextual Shadowing Effect & Curse of Overexposure
First, as shown in Fig. 3, we have collected the replay elements regarding the news (paraphrases, implications or QA pairs) during the synthetic data generation process, and assemble them properly for fine-tuning. For the aforementioned experiments, we used the single-turn conversation format as one row of the training data, where the user asks for a topic or question, and the assistant responds with a replay element.
@@
Now, we investigate what happens if we use a somewhat richer format where the news is given explicitly in the context followed by QAs as a multi-turn conversation, illustrated in Fig. 5, where the assistant responds both the news and the replay element in context. We fine-tuned models with the data in this format.
@@
Surprisingly, the prefixed context catastrophically degrades learning, as seen in Fig. 6 a,b. This effect during training is robust and consistent across Sys2-FT protocols and model scales (Fig. 41). Our interpretation is as follows: since models are able to effectively use the news via ICL, the {answer} tokens learning signal is shadowed away as it is not a surprise given the news in context. We coin such phenomenon the “Contextual Shadowing Effect” and suggest that it introduces a significant challenge in organizing data format for fine-tuning or even pre-training, as an earlier appearance of a concept in context hinders learning the concept itself and its downstream consequences. This might raise realistic concerns as scientific papers/textbooks often include a definition or fact to be applied or used later in the document, for which the contextual shadowing effect will result in poor knowledge internalization performance when naively fine-tuning with the whole document in context. We observe the contextual shadowing effect robustly across all Sys2-FT protocols (Fig. 29).
@@
Another phenomenon we observe is the curse of overexposure for in-context learning. Does fine-tuning affect the in-context learning ability of models? Surprisingly, as Fig. 6 c demonstrates, the model’s in-context learning ability sometimes gets degraded during training: as the model gets better at answering questions related to the news, the model becomes worse when the news is explicitly given in context. We coin this phenomenon the “curse of overexposure” and hypothesize that overexposure of the news during fine-tuning can harm the ICL circuit of the model. We observe this phenomenon in most training runs but not all (App. D.4). Note that this is not a unique problem with Sys2-FT, but prevalent for all FT methods. Also, we often find that training long enough recovers the ICL accuracy. With the current experiments it is hard to identify the exact factors that cause such behaviors and the “curse” merely refers to the unexpected modulation of ICL accuracy during training.
6 General Capability Preservation and Continual Learning
As model fine-tuned with the Sys2-FT method on our dataset achieve high evaluation performance on the new knowledge, an important question that remains is to what extent it will harm the preexisting capabilities of the model. In Fig. 7, we show that general knowledge and instruction following capabilities are well preserved after Self-QA System-2 Fine-tuning. In another set of experiments (Fig. 8), after fine-tuning on one subdomain of the New News dataset and merging with that LoRA (Hu et al., 2021) adapter, we continually fine-tune the model on another subdomain of the dataset and observe relatively small forgetting of the knowledge just learned. These experiments suggest that preexisting knowledge and capabilities are preserved after Self-QA System-2 fine-tuning with relatively small forgetting, as the model continually learns new information.
@@ 7 Analysis of Sys2-FT across Data Generation models
To isolate the model size effect, we conduct Self-QA FT experiments across model scales with data generated all other model scales as well. We focus on the Self-QA protocol as it is the most stable and robust across data splits and model sizes (Fig. 4). As shown in Fig. 9, we notice that the successful models lie in a rectangular region in the heatmap, which suggests that: 1) Successful integration of new knowledge requires both a sufficiently good model and high quality data, as opposed to a smooth trade off; 2) Stronger models fine-tuned with lower-quality data generated by weak models can surpass the performance of weak model itself, a phenomenon reminiscent of weak-to-strong generalization (Burns et al., 2023). Additional details on data quality generated by models of different scales and overlap with the evaluation questions can be found in Appendix C.4.
@@ 9 Discussion & Conclusion
In this paper, we introduce a new dataset New News to study whether models can integrate new knowledge consistently via fine-tuning. Inspired by memory consolidation and information replay from cognitive science, we propose the Sys2-FT method by leveraging language models’ in-context generations with news. We demonstrate that Sys2-FT is a robust method to teach and update models with new information and concepts, exemplified by the Self-QA protocol on the mathematics and coding dataset, while preserving the general capabilities after fine-tuning. Although our methods were somewhat successful and generalized to a different family of models (Llama 3.1 8B, App. C.3), integrating new knowledge into models via fine-tuning to surpass in-context learning remains a challenge, especially in domains outside mathematics and coding and in small models.
@@
CFP and ZZ identified the problem and discussed the setup. CFP and ZZ created the New News dataset. CFP ran all experiments. CFP and ZZ planned the analysis together. ZZ wrote the first draft and CFP and ZZ finalized the paper together. HT funded and supervised the research.